OpenStudy (anonymous):

Simplify by dividing - 5/8 divided by - 3/4. A. 20/24 B. - 20/24 C. - 15/32 D. 15/32 Also, every time I put a negative symbol, the whole decimal was negative, not just the dividend.

OpenStudy (michele_laino):

hint: we can write this: \[\Large \left( { - \frac{5}{8}} \right):\left( { - \frac{3}{4}} \right) = \left( { - \frac{5}{8}} \right) \times \left( { - \frac{4}{3}} \right) = ...?\]

OpenStudy (anonymous):

I'm pretty new to Openstudy, so I don't know how to do that yet. :/

OpenStudy (anonymous):

At least, not in a question.

OpenStudy (michele_laino):

please continue my computation above

OpenStudy (michele_laino):

hint: \[\left( { - \frac{5}{8}} \right):\left( { - \frac{3}{4}} \right) = \left( { - \frac{5}{8}} \right) \times \left( { - \frac{4}{3}} \right) = \frac{{\left( { - 5} \right) \times \left( { - 4} \right)}}{{8 \times 3}} = ...?\]

OpenStudy (anonymous):

Ok, I got it! Thank you so much!

OpenStudy (michele_laino):

Welcome to Open Study!!!!! :)